OpenStudy (anonymous):

1. Find the horizontal asymptote of the graph of -6x^3+5x+8/x^3+x+8 2. Simplify the rational expression. State any restrictions on the variable. p^2-4p-32/p+4 3. Simplify the rational expression. State any restrictions on the variable. q^2+11q+24/q^2-5q-24 4. Write an equation for the translation of y=-1/x that has the asymptotes x = –2 and y = 4.

OpenStudy (campbell_st):

the horizontal asymptote if easy find using polynomial division so basically it occurs at y = -6 because the numerator and denominator are of the same degree. so its really \[y = \frac{-6x^3}{x^3} = -6\] if you did the division properly there would be a remainder, which is ignored. hope this helps

OpenStudy (campbell_st):

question 2 the restrictions occur in the denominator... it can't be zero so to find the restriction solve p + 4 = 0 the answer is the restriction. so simplify the expression factorise the numerator and you'll see a binomial factor will cancel leaving a linear expression. The restriction would become a point of discontinuity rather than a asymptote

OpenStudy (campbell_st):

Question 3 similar to question 2 factorise the numerator and denominator factors will cancel. the restrictions will be the values that make the denominator zero, easy to find after factoring.

OpenStudy (campbell_st):

question 4 to get a horizontal asymptote of y =4 the numerator would need to be 4x the vertical is found by expression the asymptote x = -2 in terms of x + 2 = 0 x + 2 becomes the new denominator.

OpenStudy (anonymous):

@campbell_st so number 2 is \[p+8;p \neq 4\]

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

nope the denominator is p + 4 so solve p + 4 = 0 that is the restriction and factor \[x^2 - 4x - 32\] find the factors of -32 that add to -4, the larger factor is negative.

OpenStudy (anonymous):

thank you so much